I am having problems setting this equation up. The problem is: If two cars leave the same location at the same time, And car one is traveling 20 MPH faster than car two in opposite directions, and they are 168 miles apart 1 1/2 hours later...how fast is each one going? Your help is GREATLY needed. Thank you!
Found 2 solutions by khwang, usyim88hk:
Answer by khwang(438)

Assume the speed of car one is x mph,
then the speed of car two is x-2 mph.
When they travel in the opposite direction,
distance = sum of two speed * Time,
So, 168 = (x+x-2)* (3/2) = 3x -3,
or 3x = 171, x = 57 and x-2 = 55.
Answer: speed of car 1 is 57 MPH,
speed of car 2 is 55 MPH

Answer by usyim88hk(157)

Let x be the slower car
x = slower car
x+20 = faster car
168/1.5 = 112 = how much they are apart in one hour
x+x+20 = 112
2x = 92
x = 46
So the speed of the slower car is 46mph, the speed of the faster car is 66mph